Two angles are supplementary and one angle measures 30 less than twice the other. What is the measure of the LARGER angle? (Enter a number)
Answer by checkley77(12569)

X+Y=180 OR X=180-Y
X=2Y-30
2Y-30=180-Y
2Y+Y=180+30
3Y=210
Y=210/3
Y=70 DEGREES.
X=2*70-30
X=140-30
X=110 ANS.
PROOF:
70+110=180
180=180